OpenStudy (anonymous):

Can someone help me. 80% of what number of salads is 20?

OpenStudy (anonymous):

80% =80/100=.8 80% of a number is the same as .8*x which is equal to 20 .8*x=20 solve for x

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

sorry it took me so long too respond back.

OpenStudy (anonymous):

is it 25

OpenStudy (anonymous):

yup